OpenStudy (anonymous):

If the amount of air in a spherical balloon is 288 pi cu. in., what is the radius if half the air is let out? The answer must be in cube root form.

OpenStudy (btaylor):

\[V = \frac{4}{3}\pi r^{3}\] so solve for r if V=288pi/2=144pi \[144 \pi = \frac{4}{3} \pi r^3\]\[144 = \frac{4}{3}r^{3}\]\[108=r^{3}\]\[r=\sqrt[3]{108}\]

OpenStudy (anonymous):

You just saved me!! I would give you 50 medals if I could! Thanks <3

OpenStudy (btaylor):

no problem.